Question 4. (15 points) An individual who has automobile insurance from a certain company is randomly selected. Let Y be the number of moving violations for which the individual was cited during the last 3 years. The pmf of Y is 0 2 一 p(y) 0.25 0.10 0.05 (1). Find a, such that p(y) is a p.m.f (2). Write out the c.d.fof Y completely. (3), suppose an individual with Y violations incurs a surcharge of $(T+100r)·Calculate the expected amount of the surcharge.
Answer #1

(a)

a is got by noting that the Total Probability = 1

Thus,

a + 0.25 + 0.10 + 0.05 = 1

So,

a = 1 - 0.40 = 0.60

Thus, pdf of Y is given by:

Y P(Y)
0 0.6
1 0.25
2 0.10
3 0.05

The cdf of Y is given by:

F(Y) = 0 for y <0

= 0.6 for 0 \leq y < 1

= 0.85 for 1 \leq y < 2

= 0.95 for 2 \leq y < 3

= 1 for y \geq 3

(c)

y p yp y2 p
0 0.60 0 0
1 0.25 0.25 0.25
2 0.10 0.20 0.40
3 0.05 0.15 0.45
Total 0.60 1.10

So,

E(Y) = 0.60

E(Y2) = 1.10

So,

E(Y + 100 Y2) = E(Y) + 100 E(Y2)

=0.60 + (100 X 1.10) = 110.60

So,

Answer is:

110.60